The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $82.80 on December 31, 20Y2.

Marshall Inc.

Comparative Retained Earnings Statement

For the Years Ended December 31, 20Y2 and 20Y1

1

20Y2

20Y1

2

Retained earnings, January 1

$3,692,000.00

$3,262,000.00

3

Net income

610,000.00

540,000.00

4

Total

$4,302,000.00

$3,802,000.00

5

Dividends:

6

On preferred stock

$10,000.00

$10,000.00

7

On common stock

100,000.00

100,000.00

8

Total dividends

$110,000.00

$110,000.00

9

Retained earnings, December 31

$4,192,000.00

$3,692,000.00

Marshall Inc.

Comparative Income Statement

For the Years Ended December 31, 20Y2 and 20Y1

1

20Y2

20Y1

2

Sales

$10,860,000.00

$10,000,000.00

3

Cost of goods sold

6,000,000.00

5,460,000.00

4

Gross profit

$4,860,000.00

$4,540,000.00

5

Selling expenses

$2,160,000.00

$2,000,000.00

6

Administrative expenses

1,627,500.00

1,500,000.00

7

Total operating expenses

$3,787,500.00

$3,500,000.00

8

Income from operations

$1,072,500.00

$1,040,000.00

9

Other revenue

99,500.00

20,000.00

10

$1,172,000.00

$1,060,000.00

11

Other expense (interest)

132,000.00

120,000.00

12

Income before income tax

$1,040,000.00

$940,000.00

13

Income tax expense

430,000.00

400,000.00

14

Net income

$610,000.00

$540,000.00

Marshall Inc.

Comparative Balance Sheet

December 31, 20Y2 and 20Y1

1

20Y2

20Y1

2

Assets

3

Current assets:

4

Cash

$1,050,000.00

$950,000.00

5

Marketable securities

301,000.00

420,000.00

6

Accounts receivable (net)

584,000.00

500,000.00

7

Inventories

430,000.00

380,000.00

8

Prepaid expenses

109,000.00

20,000.00

9

Total current assets

$2,474,000.00

$2,270,000.00

10

Long-term investments

800,000.00

800,000.00

11

Property, plant, and equipment (net)

5,750,000.00

5,184,000.00

12

Total assets

$9,024,000.00

$8,254,000.00

13

Liabilities

14

Current liabilities

$882,000.00

$812,000.00

15

Long-term liabilities:

16

Mortgage note payable, 6%,

$200,000.00

$0.00

17

Bonds payable, 4%,

3,000,000.00

3,000,000.00

18

Total long-term liabilities

$3,200,000.00

$3,000,000.00

19

Total liabilities

$4,082,000.00

$3,812,000.00

20

Stockholders’ Equity

21

Preferred 4% stock, $5 par

$250,000.00

$250,000.00

22

Common stock, $5 par

500,000.00

500,000.00

23

Retained earnings

4,192,000.00

3,692,000.00

24

Total stockholders’ equity

$4,942,000.00

$4,442,000.00

25

Total liabilities and stockholders’ equity

$9,024,000.00

$8,254,000.00

5. Number of days' sales in receivables

7. Number of days’ sales in inventory

12. Return on total assets
Answer #1

Note : I provided three answers and answers are rounded so please take as per your requirements. Because round off factor is not provided in the problem.

5.Days sales in account receivables = average receivable /sales *365

=$542,000/10,860,000*365

=18.21639042 days without rounded

=18.22 days( if rounded to two decimals)

=18 days if rounded to near days

7. No of days sales in inventory = average inventory /cost of goods sold *365

=405000/6000,000*365

=24.6375 days (without rounded)

=24.64 days ( rounded to two decimals)

24 days (rounded to near days)

12. Return on assets = net income /average assets

=$610,000/$8,639,000 *100

=7.061002% without rounded

=7.06 if rounded to two decimals

=7% if rounded to near percent

Notes

Average receivable =430,000+380,000/2=405000

Average inventory = 584000+500,000/2=542,000

Average assets =9024000+8254000/2=8,639,000